# Blockchain Amplification Attack

### v2

@@ 12 Ethics
Following Tang et al. [49] on ethics of measuring blockchain P2P networks, our work consists of passive measurements (i.e., not submitting transactions). We do collect the Ethereum node IP addresses, but this is already publicly available information. As a part of responsible disclosure, we contacted the services that we discovered are running modified nodes and could be victims of this attack. One of the services appears to have fixed the vulnerability.
Acknowledgments
@@ Appendix A Blockchain lookup
We systematically examine all transactions observed in Flashbots’ dataset to ascertain their eventual inclusion on the blockchain, differentiating them from dropped transactions–those that enter the txpool but fail to be included on the blockchain. Our analysis involves comparing the blockchain information up to the future seven days from the observation timestamp. There is a potential for mislabeling a transaction if it takes longer than seven days to be a part of the blockchain. To substantiate the adequacy of the 7-day lookup window, we conducted an experiment to check the variation in our results (i.e., optimizing the extent of blockchain lookup required to define dropped transactions). Utilizing 100 days of data starting from September 1st, we compute the number of dropped transactions for each day, calibrating the length of the blockchain lookup by extending the reference point into the future.
@@
Figure 12: Number of dropped transactions.
@@ Appendix B Erigon propagation policy
This supplementary section illustrates the change in Erigon’s transaction propagation policy. Erigon had adopted “square root policy” just like Geth until v2.49.0. Refer to SendMessageToRandomPeers used in BroadcastPooledTxs. However, there have been two major changes to this initial implementation.
On August 23rd, 2023 (#8030), Erigon stopped the square root policy and started to broadcast (0x02) to every peer the node connects to. The intention is to propagate the block to every node, but the change in SendMessageToRandomPeers function affects the transaction propagation simultaneously (after version v2.49.0). Next, on Dec 4th, 2023 (#8271), Erigon decided to broadcast (0x02) to a constant number of 3 peers and announce (0x08) to 6 peers (after v2.55.0). The intention is to reduce the burden of outgoing traffic. eriogn’s analysis [14] illustrates that outbound traffic reduces from 5.5-6.5 MiB/s to 3-3.5 MiB/s. Our reconstruction method (§6.1) does not apply after v2.49.0 because an Erigon node 1) does not differentiate between 0x02 and 0x08 until v2.55.0, and 2) does not announce to all peers after v2.55.0.
@@ Appendix C Manual analysis of the empirical attack
We manually look at some attackers’ addresses and list two possible motivations for the attack. The first interpretation is to disrupt the service or the network. If the account sends many transactions with insufficient addresses or past nonce, the attacker has no intention of making transactions on-chain, but solely to cause a disturbance to the service/network (i.e., Amplification attack).
The second interpretation is to increase the chance of transaction inclusion. The intention is different from our attack model. We look into two accounts (0x3d9e…, 0x443d…) that slightly alter the gas limit, and send thousands of invariant transactions in one block slot. Those accounts appear to be the arbitrage MEV bots. By sending thousands of transactions at once, the transactions may reach out to the nodes faster than others or fill up the txpool to prevent other transactions from being executed. This behavior is partly discussed in another study by Zhou et al [57]. While many of their on-chain transactions get reverted in the middle of the execution, there are some successful complete arbitrage transactions. One transaction (0x3d1d…) has made 1.42 Ether by taking an arbitrage between Uni Swap and Sushi Swap, which would be 2,909 USD at the time of execution. As long as the cost (i.e., the amount of gas fee that the MEV bot pays for on-chain transactions and the computational cost of generating invalid transactions) does not exceed the profit of successful arbitrage, the bot has an incentive to spam the network.
@@ Appendix D MLE estimate of peer connections $\hat{\theta}_{i}$ and $\hat{x}_{i}$
Figure 13: The relationship among $m$, $x_{i}$, and the amount of error.
@@
$=-\frac{m_{2}}{\theta^{2}}+\frac{m_{8}}{(1-\theta^{2}}\quad\text{Using $\hat{\theta}_{i}=\frac{m_{2}}{m_{2}+m_{8}}$}$
@@
The variance is the negative inverse of the 2nd derivative; thus $Var[\hat{\theta}_{i}]=\frac{\hat{\theta}_{i}(1-\hat{\theta}_{i})}{m}$. It is at most (bounded by) $\frac{1}{2\sqrt{m}}$ when $\hat{\theta}=\frac{1}{2}$ (maximum). Within two standard deviation (std) $\sigma_{\hat{\theta}_{i}}$ (which covers around 95%), a confidence interval (CI) of $\hat{\theta}_{i}$ is $\hat{\theta}_{i}\pm 2\sigma_{\hat{\theta}_{i}}=\hat{\theta}_{i}+\frac{1}{\sqrt{m}}$. However, after the reconstruction of $\hat{x}_{i}$ using Eqn.(4), the variance of $\hat{x}_{i}$ does not appear to have a closed-form solution, so we use heuristics to determine the credibility of our estimate $\hat{x}_{i}$. In particular, since the variance of $\hat{x}_{i}$ would expand more quickly when $\hat{\theta}_{i}$ is small, we exclude the uncertain estimates based on the level of $\hat{\theta}_{i}$ and $Var[\hat{\theta}_{i}]$. We calculate $\hat{x}_{i}$ from two $\hat{\theta}$s: 1) $\hat{\theta}_{i}$, and 2) $\hat{\theta}_{i}+(1/\sqrt{m})$ (two std apart), and refer to the difference of the two points as an “error,” $\varepsilon$. If the error is more than 10 (i.e., the estimated number of peers $\hat{x}_{i}$ likely to deviate more than 10 peers), we exclude the estimate.
$\varepsilon=\underbrace{\left(\frac{1}{\hat{\theta}_{i}}\right)^{2}}_{\theta_{i}=\hat{\theta}_{i}}-\underbrace{\left(\frac{1}{\hat{\theta}_{i}+(1/\sqrt{m})}\right)^{2}}_{\theta_{i}=\hat{\theta}_{i}+(1/\sqrt{m})}=x_{i}-\left(\frac{1}{\sqrt{x}}+\frac{1}{\sqrt{m}}\right)^{-2}\ .$ (5)
We next look at the relationship between $x$, $m$, and the error $\varepsilon$. Figure 13 (left/middle) illustrates the change in the level of error by increasing $m$ and $x_{i}$. When the number of observations $m$ increases, the error marginally decreases. When the peer connection $x$ increases, the error slowly increases. Figure 13 (right) shows the relationship between $x$ and $m$ while fixing the amount of error. A much larger number of samples is necessary for the node with many connections to obtain a reliable estimate.
@@ Appendix E Customized nodes
Figure 14: Number active peer connections in our nodes’ connections. The $y$-axis does not start at 0.
Figure 15: Number active customized peer connections in our nodes’ connections.
@@
We identify many clusters of customized nodes. For example, the largest cluster has 250 nodes, and all reside in the same ISP (data center) in one (US) city. There are also two clusters (200+ different IDs) that appear to be operated by the same entity, but those nodes are spread out across different ISPs and 19 cities. We infer that they may belong to one relay service that attempts to reduce the latency by deploying nodes globally. Figure 14 illustrates the number of active connections in our two deployed nodes (based on the timestamp of connection/disconnection for each peer), which shows the stability of our nodes over time. Figure 15 depicts the number of customized nodes; around 100 (10%) are customized nodes.
@@ Appendix F Software client, ISP
We investigate software client market share to validate our choice of transaction propagation policies in §4.3. To exclude unstable/non-functioning nodes, we produce statistics from nodes that maintain peer connections to our two nodes. Our reported statistics are thus different from those at websites such as Etherscan [18], which include all peers from other chains in the P2P network. We observe a software market share that remains stable over time. Geth, Erigon, and Nethermind have 75%, 20%, and 5%, respectively. We can thus focus on Geth and Erigon in our model.
@@ Appendix G Attack scenario
We aim to estimate the highest possible traffic costs that modified nodes would incur when an attacker saturates their outgoing traffic with its invalid transactions. The attacker can utilize 400 Ethereum blockchain accounts (as shown in Figure 10) to evict existing transactions in the txpool and keep refreshing it with the new invalid ones, causing the modified nodes to continuously forward invalid transactions to their neighbors. The attacker can also send transactions from multiple nodes with different IPs, and distribute the attack to multiple modified nodes to scale up the attack until the bandwidth limit (i.e., Distributed EDoS attack). This helps attackers circumvent any network management tools deployed by modified nodes (e.g., anomaly detection—one of the solutions to classical EDoS attacks). We partly reference Kumar’s work [33] to calculate the total link capacity of intermediate nodes (i.e., modified nodes in our case). Table IV summarizes the assumptions and the final cost of modified nodes. We use the bandwidth of AWS EC2 instances for attacker/modified nodes and the minimum bandwidth requirement by Geth for the regular nodes. [^17] https://geth.ethereum.org/docs/getting-started/hardware-requirements In our case, the scale of the attack (the amount of waste) is mainly constrained by the number/bandwidth of modified/regular nodes since the attacker’s traffic gets easily amplified by TAF. Based on our calculation, each modified node could spend at most 88,904 USD, or roughly 8M USD per month in aggregate if all the modified nodes are from the same entity.
Alternatively, the attacker might opt to perform this attack moderately over an extended duration to evade detection, rather than maximizing traffic volume over a brief period. Previous literature [5] highlights the potential “stealthy nature” of the EDoS attack (compared to DoS) where the attacker carefully chooses the level of traffic under the detection threshold set by the victim. It might be more economically effective to control the level of the attack given the marginal decrease in EAF (Figure 9).
TABLE IV: One attack scenario ($\mathcal{M}$: modified node, $\mathcal{R}$: regular node) with link capacity constraints.
@@ Appendix H Analysis of validation simulation
Ethereum uses the MPT to maintain the blockchain state, which is a tree structure where a leaf node stores the value of persistent data (e.g., account balance), and all intermediate nodes in the path from the root node to the leaf node correspond to the key of the data’s value (e.g., account address). Fetching account nonces also requires multiple lookups in the tree. This traversal process is time-consuming for several reasons [26]. First, due to the large amount of state, the tree can become deep, increasing the number of nodes that must be traversed to find a particular piece of data. Second, disk access plays a significant role because the MPT structure often resides on disk. Slow disk accesses occur when the nodes required for a lookup are not in memory and must be fetched from disk.
@@
Figure 16: Profiling execution time for transaction validation — analysis of $136,437$ Transactions from $1,000$ consecutive blocks starting at block $18,140,000$ using Golang’s “pprof.”

### v3

@@ 9.3. Ethics
Following Tang et al. (Tang et al., 2023b) on ethics of measuring blockchain P2P networks, our work consists of passive measurements (i.e., not submitting transactions ourselves). We do collect the Ethereum node IP addresses, but this is already publicly available information. In an effort to perform responsible disclosure, we contacted the services that we discovered are running modified nodes and could be victims of this attack. At least one of those services appears to have fixed the vulnerability.
10. Related work
We discuss related work in P2P network measurements and network security.
10.1. P2P network measurements
Most measurements on blockchain P2P networks focus on the network topology such as network size (Kim et al., 2018; Neudecker et al., 2016), influential nodes (Li et al., 2021a; Maeng et al., 2021), connection stabilities/longevities (Neudecker, 2019; Daniel et al., 2019), and latency between peers (Decker and Wattenhofer, 2013; Gencer et al., 2018; Tang et al., 2023a; Wahrstätter et al., 2024). Some websites summarize those statistics in real-time for various networks: Bitcoin (bitnodes.io, 2023; coin.dance, 2023; Systems and KASTEL, 2023); Ethereum (Etherscan.io, 2023; Ethernodes.org, 2023; Mazza, 2023); Monero (MoneroHash.com, 2023).
In particular, there has been extensive work on identifying active peer connections to further understand the propagation flow over the network. The techniques developed include: utilizing latency between peers (Neudecker et al., 2016; Daniel et al., 2019), isolation property and orphan transactions (Delgado-Segura et al., 2019), connection freshness (Cao et al., 2020), and gas fee change (Li et al., 2021a; Zhao et al., 2024). However, these works are either 1) chain-specific and do not apply to Ethereum, and/or 2) involve active measurement (e.g., submitting transactions), which is expensive and could disrupt the main network. Our proposed method in §6.1 works on the Ethereum main network 1) without intervening in ongoing activities and 2) without incurring any additional costs apart from collecting data. Our method only considers the number of connections and does not determine whether an active edge exists between two specific nodes, making it harder to compare its accuracy with previous proposals.
Another body of measurement research focuses on characterizing nodes, e.g., centralized entities such as mining or relay nodes (Miller et al., 2015; Daniel et al., 2019; Li et al., 2021a, b), AS/cloud services (Feld et al., 2014; Gencer et al., 2018; Neudecker et al., 2016), geo-distributions (Neudecker, 2019; Kim et al., 2018; Daniel et al., 2019; Maeng et al., 2021; Cao et al., 2020; Cortes-Goicoechea et al., 2023), and peer technical specifications (i.e., bandwidth, CPUs) (Tang et al., 2023a; Cortes-Goicoechea et al., 2023). Closest to our work is the literature about the client (software) types – for Bitcoin (Neudecker, 2019), Ethereum (Kim et al., 2018; Cortes-Goicoechea et al., 2023; Grandjean et al., 2023), and Zcash (Daniel et al., 2019). Those works mostly focus on the diversity of software and/or versions, with relatively scarce attention on the implementation and performance. To the best of our knowledge, there has not been any exploration of customized clients, a gap our study aims to bridge.
10.2. P2P network security: DoS and spam
Due to the rise of centralized entities—mining pools, exchanges, and relay nodes, Denial of Service (DoS) attacks on those services can have a disastrous impact on the overall network (Vasek et al., 2014; Gervais et al., 2015). The most relevant work is by Li et al (Li et al., 2021b) (and its extension by Yaish et al (Yaish et al., 2023)), which sends transactions with a future nonce to evict legitimate transactions in the txpool. Further, Wang et al (Wang et al., 2023) develop a fuzzer to automatically discover txpool vulnerabilities. Heo et al (Heo et al., 2023) illustrate that attackers can delay honest transaction execution by sending invalid transactions. Zhou et al (Zhou et al., 2021a) document the spamming behavior of MEV bots as a computational overhead to the network. Our paper describes a new blockchain DoS attack, exploiting the lack of transaction validations and empirically derives the amplification factors. We also reference some metrics from non-blockchain DoS literature such as amplified DoS and EDoS attacks from Kumar (Kumar, 2007) and Wang et al. (Wang et al., 2016), respectively.
11. Conclusion
Latency reduction in blockchain P2P networks such as Ethereum can create economic value as a form of extractable value (MEV/BEV); as a result, some entities modify software clients (“modified nodes”) to shorten the transaction validation process. Unfortunately, these modifications enable a Blockchain Amplification Attack. We formalize this attack, illustrate its practicality from similar attack instances observed in the wild, and empirically measure and simulate its impact. Specifically, we show that every byte a single attacker sends can result in 3,600 bytes traveling on the network; likewise, every dollar the attacker spends on the attack can result in \$13,800 worth of economic loss. We also quantify the benefits of skipping validation based on our simulations and measurements, highlighting the economic trade-off centralized entities face, when deciding whether to accept the risk of attack. We conclude by noting that a lack of transaction validation not only increases traffic costs on the modified nodes, but also introduces invalid transactions and degrades the user experience. Our models provide a foundation for explaining how invalid transactions propagates through the network and pose a threat to the overall blockchain ecosystem.
Acknowledgements.
@@ Appendix A Blockchain lookup
This section validates our definition of dropped transactions. We systematically examine all transactions observed in Flashbots’ dataset to ascertain their eventual inclusion on the blockchain, differentiating them from dropped transactions—those that enter the txpool but fail to be included on the blockchain. Our analysis involves comparing the blockchain information up to the future seven days from the observation timestamp. There is a potential for mislabeling a transaction if it takes longer than seven days to be a part of the blockchain. To substantiate the adequacy of the 7-day lookup window, we experiment to check the variation in our results (i.e., optimizing the extent of blockchain lookup required to define dropped transactions). Utilizing 100 days of data starting from September 1st, we compute the number of dropped transactions for each day, calibrating the length of the blockchain lookup by extending the reference point into the future.
@@
Figure 9. Number of dropped transactions.
@@ Appendix B Erigon propagation policy
This section illustrates the change in Erigon’s transaction propagation policy. Erigon had adopted “square root policy” just like Geth until v2.49.0. Refer to SendMessageToRandomPeers used in BroadcastPooledTxs. However, there have been two major changes to this initial implementation.
On August 23rd, 2023 (#8030), Erigon stopped the square root policy and started to broadcast (0x02) to every peer the node connects to. The intention is to propagate the block to every node, but the change in SendMessageToRandomPeers function affects the transaction propagation simultaneously (after version v2.49.0). Next, on Dec 4th, 2023 (#8271), Erigon decided to broadcast (0x02) to a constant number of 3 peers and announce (0x08) to 6 peers (after v2.55.0). The intention is to reduce the burden of outgoing traffic. eriogn’s analysis (erigon, 2024) illustrates that outbound traffic reduces from 5.5-6.5 MiB/s to 3-3.5 MiB/s. Our reconstruction method (§6.1) does not apply after v2.49.0 because an Erigon node 1) does not differentiate between 0x02 and 0x08 until v2.55.0, and 2) does not announce to all peers after v2.55.0.
@@ Appendix C Manual analysis of the empirical attack
We manually look at some attackers’ addresses and list two possible motivations for the attack. The first interpretation is to disrupt the service or the network. If the account sends many transactions with insufficient addresses or past nonce, the attacker has no intention of making transactions on-chain, but solely to cause a disturbance to the service/network (i.e., an amplification attack). Figure 10 summarizes the number of attacks over time (on an hourly basis). There are more than 60 instances of attacks on November 26th between 7–8 AM (UTC). Many of them appear to send transactions from accounts with zero balances. We find similar cases on September 12th, 5 PM, October 7th, 5 PM, and November 20th, 9 PM. The number of attacks does not appear to be correlated with the Ether price fluctuation.
Figure 10. Number of attacks observed per hour, over time (September 1st, 2023–January 11th, 2024).
The second interpretation is to increase the chance of transaction inclusion. The intention is different from our attack model. We look into two accounts (0x3d9e…, 0x443d…) that slightly alter the gas limit, and send thousands of invariant transactions in one block slot. Those accounts appear to be the arbitrage MEV bots. By sending thousands of transactions at once, the transactions may reach out to the nodes faster than others or fill up the txpool to prevent other bots’ transactions from being executed. This behavior is partly discussed in another study by Zhou et al (Zhou et al., 2021a). While many of their on-chain transactions get reverted in the middle of the execution, there are some successful complete arbitrage transactions. One transaction (0x3d1d…) has made 1.42 Ether by taking an arbitrage between Uni Swap and Sushi Swap, which would be 2,909 USD at the time of execution. As long as the cost (i.e., the amount of gas fee that the MEV bot pays for on-chain transactions and the computational cost of generating invalid transactions) does not exceed the profit of successful arbitrage, the bot has an incentive to spam the network.
@@ Appendix D MLE estimate of peer connections $\hat{\theta}_{i}$ and $\hat{x}_{i}$
Figure 11. The relationship among $m$, $x_{i}$, and the amount of error.
@@
$=-\frac{m_{2}}{\theta^{2}}+\frac{m_{8}}{(1-\theta)^{2}}\quad\text{Using $\hat{\theta}_{i}=\frac{m_{2}}{m_{2}+m_{8}}$}$
@@
The variance is the negative inverse of the 2nd derivative; thus $Var[\hat{\theta}_{i}]=\frac{\hat{\theta}_{i}(1-\hat{\theta}_{i})}{m}$. It is at most (bounded by) $\frac{1}{2\sqrt{m}}$ when $\hat{\theta}=\frac{1}{2}$ (maximum). Within two standard deviation (std) $\sigma_{\hat{\theta}_{i}}$ (which covers around 95%), a confidence interval (CI) of $\hat{\theta}_{i}$ is $\hat{\theta}_{i}\pm 2\sigma_{\hat{\theta}_{i}}=\hat{\theta}_{i}+\frac{1}{\sqrt{m}}$. However, after the reconstruction of $\hat{x}_{i}$ using Eqn.(3), the variance of $\hat{x}_{i}$ does not appear to have a closed-form solution, so we use heuristics to determine the credibility of our estimate $\hat{x}_{i}$. In particular, since the variance of $\hat{x}_{i}$ would expand more quickly when $\hat{\theta}_{i}$ is small, we exclude the uncertain estimates based on the level of $\hat{\theta}_{i}$ and $Var[\hat{\theta}_{i}]$. We calculate $\hat{x}_{i}$ from two $\hat{\theta}$s: 1) $\hat{\theta}_{i}$, and 2) $\hat{\theta}_{i}+(1/\sqrt{m})$ (two std apart), and refer to the difference of the two points as an “error,” $\varepsilon$. If the error is more than 10 (i.e., the estimated number of peers $\hat{x}_{i}$ is likely to deviate more than 10 peers), we exclude the estimate.
(4) $\varepsilon=\underbrace{\left(\frac{1}{\hat{\theta}_{i}}\right)^{2}}_{\theta_{i}=\hat{\theta}_{i}}-\underbrace{\left(\frac{1}{\hat{\theta}_{i}+(1/\sqrt{m})}\right)^{2}}_{\theta_{i}=\hat{\theta}_{i}+(1/\sqrt{m})}=x_{i}-\left(\frac{1}{\sqrt{x}}+\frac{1}{\sqrt{m}}\right)^{-2}\ .$
We next look at the relationship between $x$, $m$, and the error $\varepsilon$. Figure 11 (left/middle) illustrates the change in the level of error by increasing $m$ and $x$. When the number of observations $m$ increases, the error marginally decreases. When the peer connection $x$ increases, the error slowly increases. Figure 11 (right) shows the relationship between $x$ and $m$ while fixing the amount of error. A much larger number of samples is necessary for the node with many connections to obtain a reliable estimate.
@@ Appendix E Customized nodes
Figure 12. Number active peer connections in our nodes’ connections. The $y$-axis does not start at 0.
Figure 13. Number active customized peer connections in our nodes’ connections.
@@
We identify many clusters of customized nodes. For example, the largest cluster has 250 nodes, and all reside in the same ISP (data center) in one (US) city. There are also two clusters (200+ different IDs) that appear to be operated by the same entity, but those nodes are spread out across multiple ISPs and 19 cities. We infer that they may belong to one relay service that attempts to reduce the latency by deploying nodes globally. Figure 12 illustrates the number of active connections in our two deployed nodes (based on the timestamp of connection/disconnection for each peer), which shows the stability of our nodes over time. Figure 13 depicts the number of customized nodes; around 100 (10%) are customized nodes.
@@ Appendix F Software client, ISP
We investigate software client market share to validate our choice of transaction propagation policies in §4.3. To exclude unstable/non-functioning nodes, we produce statistics from nodes that maintain peer connections to our two nodes. Our reported statistics are thus different from those at websites such as Etherscan (Etherscan.io, 2023), which include all peers from other chains in the P2P network. We observe a software market share that remains stable over time. Geth, Erigon, and Nethermind have 75%, 20%, and 5%, respectively. We can thus focus on Geth and Erigon in our model.
@@ Appendix G Attack scenario
We aim to estimate the highest possible traffic costs that modified nodes would incur when an attacker saturates their outgoing traffic with its invalid transactions. The attacker can utilize at most 400 Ethereum blockchain accounts (as shown in Figure 8) to evict existing transactions in the txpool and keep refreshing it with the new invalid ones, causing the modified nodes to continuously forward invalid transactions to their neighbors. The attacker can also send transactions from multiple nodes with different IPs, and distribute the attack to multiple modified nodes to scale up the attack until the bandwidth limit (i.e., Distributed EDoS attack). This helps attackers circumvent any network management tools deployed by modified nodes (e.g., anomaly detection—one of the solutions to classical EDoS attacks). We partly reference Kumar’s work (Kumar, 2007) to calculate the total link capacity of intermediate nodes (i.e., modified nodes in our case). Table 3 summarizes the assumptions and the final cost of modified nodes. We use the bandwidth of AWS EC2 instances for attacker/modified nodes and the minimum bandwidth requirement by Geth for the regular nodes. [^14] https://geth.ethereum.org/docs/getting-started/hardware-requirements In our case, the scale of the attack (the amount of waste) is mainly constrained by the number/bandwidth of modified/regular nodes since the attacker’s traffic gets easily amplified by TAF. Based on our calculation, each modified node could spend at most 88,904 USD, or roughly 8M USD per month in aggregate if all the modified nodes are from the same entity.
@@ 32.4 TB/month 191,484 TB/month
Table 3. One attack scenario ($\mathcal{M}$: modified node, $\mathcal{R}$: regular node) with link capacity constraints.
Alternatively, the attacker might opt to perform this attack moderately over an extended duration to evade detection, rather than maximizing traffic volume over a brief period. Previous literature (Chowdhury et al., 2017) highlights the potential “stealthy nature” of the EDoS attack (compared to DoS) where the attacker carefully chooses the level of traffic under the detection threshold set by the victim.
This strategy is indeed effective under the AWS’s pay-as-you-go pricing model. AWS charges 90 USD/TB until 10TB, 85 USD/TB for the next 40TB, 70 USD/TB for the next 100TB, and 50 USD greater than 150TB. Figure 14 illustrates the decrease in EAF ($y$-axis) based on the amount of outgoing traffic ($x$-axis) due to AWS’s dynamic pricing. The attacker’s economic benefit marginally decreases over the scale of the attack. It might be more economically effective to control the level of the attack.
Figure 14. EAF based on the traffic usage of a modified node.
@@ Appendix H Analysis of validation simulation
Ethereum uses the MPT to maintain the blockchain state, which is a tree structure where a leaf node stores the value of persistent data (e.g., account balance), and all intermediate nodes in the path from the root node to the leaf node correspond to the key of the data’s value (e.g., account address). Fetching account nonces also requires multiple lookups in the tree. This traversal process is time-consuming for several reasons (He et al., 2024). First, due to the large amount of state, the tree can become deep, increasing the number of nodes that must be traversed to find a particular piece of data. Second, disk access plays a significant role because the MPT structure often resides on disk. Slow disk accesses occur when the nodes required for a lookup are not in memory and must be fetched from disk.
@@
Figure 15. Profiling execution time for transaction validation — analysis of $136,437$ Transactions from $1,000$ consecutive blocks starting at block $18,140,000$ using Golang’s “pprof.”